Define each of the following three (3) terms that are used for a bond:

(a) coupon rate;

(b) current yield; and

(c) yield to maturity (YTM).

Then, tell me how these 3 different rates would relate to each other (ie. rank them from the highest to the lowest) if the bond were selling at a premium.
Answer #1

Coupon Rate - A bond carry a specific rate at the time of issue of bond is known as coupon rate.In other words, the rate of return already given at the time of issue of bond, like 7.5% $1,000 Bond, the interset it'll earn is 75$.

Current Yield- Here, in Current Yield, the current price of the bond is taken into consideration rather than its face value and compared with the annual income from the bond, to determine the yield.

Its formula is = \frac{Annual Income from Bond}{Purchase Price}

Yield to Maturity(YTM)- YTM is the potential rate of earning from the bond, if the bond is held till its maturity.

YTM= \frac{Coupon Payment + \frac{Face Value- Price}{Number of Years}}{\frac{Face Value+Price}{2}}

In case, the bond is selling at a premium i.e the value is above the face value, then the Coupon Rate is higher than its YTM and Current price.

Ranking it from highest to lowest-

1. Coupon Rate

2. Current Yield

3. YTM